convert between degress fahrenheit and degrees celsius using the literal equation given:
c=(5/9)(f-32) find the celsius when f=100

Answer by ReadingBoosters(3246)

c = 5/9(f-32) where f=100
c = 5/9(100-32)
c = 5/9(68)
c = 340/9
c = 37.78 degrees
............
Delighted to help.
-Reading Boosters
Wanting for other what we want for ourselves.
www.MyHomeworkAnswers.com